OpenStudy (anonymous):

A ball is thrown from a height of 156 feet with an initial downward velocity of 8/fts. The ball's height h (in feet) after t seconds is given by the following. h=156−8t-6t^2 How long after the ball is thrown does it hit the ground?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

h=156−8t-6t^2 0 = 156−8t-6t^2

OpenStudy (anonymous):

@Directrix can't you solve it using quadratic formula?

Directrix (directrix):

I'm trying to decide if the 0 is in the right place.

OpenStudy (anonymous):

4.47 -5.80 before i round up

OpenStudy (anonymous):

i used quadrtic formula

OpenStudy (anonymous):

\[a x^{2}+bx+c=0\] So a would be 6 b would be 8 c would be -156 And quadratic formula would be \[\frac{ -b \pm \sqrt{b ^{2}-ac} }{ 2a }\] Then plug and chug

OpenStudy (anonymous):

I would set it to 0

Directrix (directrix):

Let me see what I get. I think that the zero is in the right place.

OpenStudy (anonymous):

i havent inputted yet i wait for u to tell me if im right

Directrix (directrix):

>>4.47 -5.80 I got these but which one is the answer? And what about the rounding?

OpenStudy (anonymous):

i have a button for or so i have to put both

OpenStudy (anonymous):

4.48 -5.81 ??

Directrix (directrix):

Okay. Any instructions about rounding?

OpenStudy (anonymous):

Round at the very end

OpenStudy (anonymous):

nearest hundreth

OpenStudy (anonymous):

I also get 4.48

Directrix (directrix):

Wait a minute.

OpenStudy (anonymous):

im waiting

Directrix (directrix):

-5.81 and 4.48 @jaeuni Did you get that?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

incorrect

Directrix (directrix):

Oh, no.

Directrix (directrix):

Any comments given about the answer?

OpenStudy (anonymous):

no just incorrect

Directrix (directrix):

Do you have another shot or is this it?

OpenStudy (anonymous):

Did you type in both -5.81 and 4.48?

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

yes

Directrix (directrix):

Let me see who is online and might know what's wrong with this problem's solution.

OpenStudy (anonymous):

it gives me 2 chances if i get first wrong it makes e do another question

OpenStudy (anonymous):

Isn't it just the positive value since you can't have negative seconds in real life scenarios?

OpenStudy (anonymous):

all the others accepted the neg

OpenStudy (anonymous):

Interesting...

OpenStudy (anonymous):

thats what i thought

OpenStudy (anonymous):

it booted me out and gave the answer 2.88

zepdrix (zepdrix):

Correct. No negative time :) 4.48 should have been accepted D: 2.88? weirddddd

Directrix (directrix):

Thanks. @zepdrix